If
3.58 g
of
CuNO3
is dissolved in water to make a
0.670 M
solution, what is the volume of the solution
1
Expert's answer
2019-07-26T03:16:51-0400

1) Find moles of CuNO3

n=m/M

n(CuNO3) = 3.58/125.55 = 0.0285 mol


2) find volume of solution:

c=n/V

V=n/c = 0.0285/0.670 =0.0425 L = 42.5 mL